# Kermut: Composite kernel regression for protein variant effects

### v2

@@ 3.3.1 Zero-shot mean function
Kermut can be used with a constant mean function, $m(\mathbf{x})=\alpha$, where $\alpha$ is a hyperparameter optimized through the marginal likelihood. However, we posit that additional performance can be gained by using an altered mean function which operates on zero-shot fitness estimates, which are often available at relatively low cost: $m(\mathbf{x})=\alpha f_{0}(\mathbf{x})+\beta$, where $f_{0}$ is a zero-shot method evaluated on input sequence $\mathbf{x}$. We use ESM-2 [12], which yields the log-likelihood ratio between the variant and wild type residue as in [9].
@@ 3.4 Architecture considerations
While Kermut is based on relatively simple principles, its components are non-trivial in that they exploit learned features from pretrained models: ESM-2 provides protein sequence level embeddings and zero-shot scores, while ProteinMPNN provides a featurization of the local structural environments. We stress that these pretrained components can be readily replaced by other pretrained models. Models that generate (1) protein sequence embeddings, (2) structure-conditioned amino acid distributions, and (3) zero-shot scores are plentiful and such models are likely to progress further in future years. In our work, we have not sought to find the optimal combination of these. Our work should instead be seen as a framework demonstrating how such components can be meaningfully combined in a GP setting to obtain state-of-the-art results.
@@
No structure kernel $k_{\text{struct}}=0$ $-0.084$ $-0.078$ $-0.037$ $-0.066$
No sequence kernel $k_{\text{seq}}=0$ $-0.040$ $-0.047$ $-0.052$ $-0.046$
No inter-residue dist. $k_{d}=1$ $-0.050$ $-0.051$ $-0.007$ $-0.035$
No mut. prob./site comp. $k_{p}=k_{H}=1$ $-0.037$ $-0.038$ $-0.011$ $-0.028$
Const. mean $m(\mathbf{x})=\alpha$ $-0.039$ $-0.033$ $-0.012$ $-0.027$
No mut. prob. $k_{p}=1$ $-0.023$ $-0.019$ $-0.010$ $-0.017$
No site comp. $k_{H}=1$ $-0.007$ $-0.009$ $-0.004$ $-0.006$
@@ 4 Results
Our model reaches both higher Spearman correlations and lower mean squared error than competing methods and thereby achieves state-of-the-art performance in all three schemes, with the largest gains in the challenging modulo and contiguous schemes. Moreover, Kermut is significantly faster compared to deep learning methods. We provide wall-clock times for running a 5-fold CV loop for a single split-scheme for four select datasets for both Kermut and ProteinNPT in Table 4. Generating results for all three split schemes for the four datasets thus takes Kermut approximately 10 minutes while ProteinNPT takes upwards of 400 hours.
The non-parametric bootstrap standard error for each model relative to Kermut can be seen in Tables 5 and 7. In Appendix L, we provide additional results using alternate zero-shot functions. We also include the performance per functional category, showing similar results (Appendix E). The average and per-split performance for individual assays can be seen in Figures 16, 17, 18 and 19 while additional details on computation time can be seen in Section B.3.
@@ 4.1 Ablation study
To examine the impact of Kermut’s components, we conduct an ablation study, where we ablate its two main kernels from Equation 4 – the structure and sequence kernels – as well as the structure kernel’s subcomponents. We similarly investigate the importance of the zero-shot mean function. The ablation study is carried out on all split schemes on a subset of 174 datasets. The difference between the ablation results and the Kermut results can be seen in Table 2, where larger values indicate large component importance. For the absolute values, see Appendix F.
As indicated by the largest drop in performance, the single most important component is the structure kernel. While removing the sequence kernel leads to comparable decreases for all three schemes, removing the structure kernel primarily leads to drops in the challenging contiguous and modulo schemes. This shows that the structure kernel is crucial for characterizing unseen sites in the wild type protein. While removing the site comparison and mutation probability kernels leads to small and medium drops in performance, we observe that removing both leads to an even larger performance drop, indicating an interplay between the two. We lastly see that the inclusion of a zero-shot mean function provides a decent performance increase.
@@ 4.2 Uncertainty quantification per mutation domain
Figure 2 shows the distributions of mean predictive variances in the six domains. In the three single mutant domains, we observe that the uncertainties increase from scheme to scheme, reflecting the difficulties of the tasks and analogously the expected performance scores (Table 1). Suprisingly, when training on both single and double mutants (1M/2M), we observe a lower uncertainty on double mutants than single mutants. For many of the multi-mutant datasets, the mutants are not uniformly sampled but often include a fixed single mutation. A possible explanation is thus that it might be more challenging to decouple the signal from a double mutation into its constituent single mutation signals. In the extrapolation setting, we observe large predictive uncertainties, as expected. One explanation of the discrepancy between the variance distributions in the multi-mutant domains might lie in the difference in target distributions between training and test sets. Figure 4 in the appendix shows the overall target distribution of assays for the 51 considered multi-mutant datasets. The single and double mutants generally belong to different modalities, where the double mutants often lead to a loss of fitness. This shows the difficulty of predicting on domains not encountered during training. For reference, we include the results for the multi-mutant domains in Table 12 in the appendix.
@@
BLAT_ECOLX 0.825 0.832 0.907 0.855 4996 286 Organismal fitness [36]
PA_I34A1 0.322 0.485 0.542 0.450 1820 716 Organismal fitness [37]
TCRG1_MOUSE 0.660 0.637 0.891 0.729 621 37 Stability [4]
OPSD_HUMAN 0.653 0.699 0.866 0.691 165 348 Expression [38]
4.3 Uncertainty calibration analysis
To clarify the relationship between model uncertainty and expected performance, we proceed with a calibration analysis. First, we perform a confidence interval-based calibration analysis [39], resulting in calibration curves which in the classification setting are known as reliability diagrams [40]. The results for each dataset are obtained via five-fold cross validation, corresponding to five separately trained models for each split scheme. We select four diverse datasets as examples (Table 3), reflecting both high and low predictive performance. The mean calibration curves can be seen in Figure 3(a). For method details and results across all datasets, see Appendix I. The mean expected calibration error (ECE) is shown in the bottom of each plot, where a value of zero indicates perfect calibration. Overall, the uncertainties appear to be well-calibrated both qualitatively from the curves and quantitatively from the ECEs. Even the smallest dataset (fourth row, $N=165$) achieves decent calibration, albeit with larger variances between folds.
While the confidence interval-based analysis suggested that the uncertainty estimates were well-calibrated with some under-confidence, the same is not as visibly clear for the error-based calibration plots. We do however see a trend of increasing error with increasing uncertainty in three of the four datasets, where the curves lie close to the diagonal (as indicated by the dashed line). The second row shows poorer calibration – particularly in the modulo and contiguous schemes. The curves however remain difficult to interpret, in part due to the errorbars on both axes. To alleviate this, we compute similar metrics for all 217 datasets across the three splits, which indicate that, though varying, most calibration curves are well-behaved (see Section I.3).
We supplement the above calibration curves with Figures 7, 8, 9 and 10 in the Appendix, which show the true values plotted against the predictions. These highlight the importance of well-calibrated uncertainties and underline their role in interpreting model predictions and their trustworthiness.
@@ 4.3.1 Comparison with ProteinNPT
Monte Carlo (MC) dropout [42] is a popular uncertainty quantification technique for deep learning models. Calibration curves for ProteinNPT with MC dropout for the same four datasets across the three split schemes can be seen in Figure 11 while figures showing the true values plotted against the predictions with uncertainties are shown in Section K.3. These indicate that employing MC dropout on a deep learning model like ProteinNPT seems to provide lower levels of calibration, providing overconfident uncertainties across assays and splits. Due to the generally low uncertainties and the resulting difference in scales, the calibration curves are often far from the diagonal. The trends in the calibration curves however show that the model errors often correlate with the uncertainties. This suggests that the model can be recalibrated to achieve decent calibration.
@@ 5 Discussion
We have shown that a carefully constructed Gaussian process is able to reach state-of-the-art performance for protein variant effect prediction while providing reasonably well-calibrated uncertainty estimates. For a majority of datasets, this is achieved orders of magnitude faster than competing methods.
While the predictive performance on the substitution benchmark is an improvement over previous methods, our proposed model has its limitations. Due to the site-comparison mechanism, our model is unable to handle insertions and deletions as it only operates on a fixed structure. Additionally, as the number of mutations increases, the assumption of a fixed structure might worsen, depending on the introduced mutations, which can affect reliability. While this limitation can be fixed by online protein structure prediction, the computational cost would increase significantly. This however applies to all structure-based methods in the context of introducing large numbers of mutations and is thus not particular to Kermut. An additional limitation is the GP’s $\mathcal{O}(N^{3})$ scaling with dataset size. While not a major obstacle in the single mutant setting, dataset sizes can quickly grow when handling multiple mutants. The last decades have however produced a substantial literature on algorithms for scaling GPs to larger datasets [43, 44, 45], which could alleviate the issue, and we therefore believe this to be a technical, rather than fundamental limitation.
@@
This work was funded in part by Innovation Fund Denmark (1044-00158A), the Novo Nordisk Synergy grant (NNF200C0063709), the Novo Nordisk Foundation through the MLSS Center (Basic Machine Learning Research in Life Science, NNF20OC0062606), and the Pioneer Centre for AI (DNRF grant number P1).
@@
Appendix A GP details
A.1 Structure kernel
@@
A.2 Sequence kernel
@@ A.3 Parametrization note
A.4 Zero-shot mean function
@@ A.5 Kernel proof
Appendix B Implementation details
B.1 General details
We build our kernel using the GPyTorch framework [49]. We assume a homoschedastic Gaussian noise model, on which we place a HalfCauchy prior [50] with scale $0.1$. We fit the hyperparameters by maximizing the exact marginal likelihood with gradient descent using the AdamW optimizer [51] with learning rate 0.1 for a 150 steps, which proved to be sufficient for convergence for a number of sampled datasets. For larger, more complex datasets, alternate learning rates and optimization steps might lead to improved results.
B.2 System details
@@ B.3 Compute time
Scaling the experiments to the full ProteinGym benchmark is however costly. We were able to train/evaluate Kermut on 205/217 datasets using an NVIDIA A40 GPU with 48GB VRAM. The remaining 12 datasets were too large to fit into GPU memory without resorting to reduced precision. For these, we trained and evaluated the model using CPU only which takes considerable time. These datasets are
  • HMDH_HUMAN_Jiang_2019
  • HSP82_YEAST_Flynn_2019
  • MSH2_HUMAN_Jia_2020
  • MTHR_HUMAN_Weile_2021
  • POLG_CXB3N_Mattenberger_2021
  • POLG_DEN26_Suphatrakul_2023
  • Q2N0S5_9HIV1_Haddox_2018
  • RDRP_I33A0_Li_2023
  • S22A1_HUMAN_Yee_2023_abundance
  • S22A1_HUMAN_Yee_2023_activity
  • SC6A4_HUMAN_Young_2021
  • SHOC2_HUMAN_Kwon_2022
B.3.1 Example wall-clock time
The wall-clock times for generating the results used for the calibration curves in Figures 3(b), 3(a) and 11 for one split scheme can be seen in Table 4 for Kermut and ProteinNPT. The experiments were carried out using identical hardware. The test system is however a shared compute cluster with sharded GPUs, so variance is expected between runs. Generating the full results for the figures thus takes Kermut approximately 10 minutes while ProteinNPT takes 400 hours. This shows the significant reduction in computational burden that Kermut allows for. Both ProteinNPT and Kermut assumes that sequence embeddings are available a priori.
Table 4: Approximate wall clock times for training and evaluating Kermut and ProteinNPT for a single split scheme, i.e., by using 5-fold cross validation. While the runtime of Kermut scales with dataset size, ProteinNPT appears to scale more strongly with sequence length due to the tri-axial attention.

### v3

@@ 3.3.1 Zero-shot mean function
Kermut can be used with a constant mean function, $m(\mathbf{x})=\alpha$, where $\alpha$ is a hyperparameter optimized through the marginal likelihood. However, we posit that additional performance can be gained by using an altered mean function which operates on zero-shot fitness estimates, which are often available at relatively low cost: $m(\mathbf{x})=\alpha f_{0}(\mathbf{x})+\beta$, where $f_{0}$ is a zero-shot method evaluated on input sequence $\mathbf{x}$. This is similar to the approach employed in [37], where Rosetta scores are used as a biophysical prior. We use ESM-2 [16], which yields the log-likelihood ratio between the variant and wild type residue as in [10]. For details, see Appendix B.
@@ 3.4 Architecture considerations
While Kermut is based on relatively simple principles, its components are non-trivial in that they exploit learned features from pretrained models: ESM-2 provides protein sequence level embeddings and zero-shot scores, while ProteinMPNN provides a featurization of the local structural environments. We stress that these pretrained components can be readily replaced by other pretrained models. Models that generate (1) protein sequence embeddings, (2) structure-conditioned amino acid distributions, and (3) zero-shot scores are plentiful and such models will progress further in future years. In our work, we have not sought to find the optimal combination of these. Our work should instead be seen as a framework demonstrating how such components can be meaningfully combined in a GP setting to obtain state-of-the-art results.
@@
No structure kernel $k_{\text{struct}}=0$ $-0.082$ $-0.078$ $-0.033$ $-0.065$
No sequence kernel $k_{\text{seq}}=0$ $-0.040$ $-0.046$ $-0.048$ $-0.045$
No inter-residue dist. $k_{d}=1$ $-0.049$ $-0.051$ $-0.004$ $-0.035$
No mut. prob./site comp. $k_{p}=k_{H}=1$ $-0.035$ $-0.037$ $-0.006$ $-0.026$
Const. mean $m(\mathbf{x})=\alpha$ $-0.036$ $-0.032$ $-0.008$ $-0.026$
No mut. prob. $k_{p}=1$ $-0.022$ $-0.019$ $-0.005$ $-0.016$
No site comp. $k_{H}=1$ $-0.006$ $-0.009$ $0.000$ $-0.006$
@@ 4 Results
Our model reaches both higher Spearman correlations and lower mean squared errors than competing methods and thereby achieves state-of-the-art performance in all three schemes, with the largest gains in the challenging modulo and contiguous settings. In Table E.2, the performance per functional category is shown, where we observe a significant performance increase in binding- and stability-related prediction tasks, likely explained by inclusion of the structure kernel. In addition to its high accuracy, Kermut is significantly faster compared to deep learning methods. We provide wall-clock times for running a 5-fold CV loop for a single split-scheme for four select datasets for both Kermut and ProteinNPT in Table C.1. Generating results for all three split schemes for the four datasets thus takes Kermut approximately 10 minutes while ProteinNPT takes upwards of 400 hours.
The non-parametric bootstrap standard error for each model relative to Kermut can be seen in Tables E.1 and E.3. In Appendix M, we provide additional results using alternate zero-shot functions. The average and per-split performance for individual assays can be seen in Figures O.1, O.2, O.3 and O.4 while additional details on computation time can be seen in Appendix C. Lastly, visualizations of the distributions of optimized hyperparameters can be seen in Appendix N.
@@ 4.1 Ablation study
To examine the impact of Kermut’s components, we conduct an ablation study, where we ablate its two main kernels from Equation 4 – the structure and sequence kernels – as well as the structure kernel’s subcomponents. We similarly investigate the importance of the zero-shot mean function. The ablation study is carried out on all split schemes on a subset of 174 datasets. The difference between the ablation results and the Kermut results can be seen in Table 2, where larger values indicate large component importance. For the absolute values, see Appendix F, where we additionally include alternative kernel formulations for both structure and sequence kernels as well as for kernel composition.
As indicated by the largest drop in performance, the single most important component is the structure kernel. While removing the sequence kernel leads to comparable decreases for all three schemes, removing the structure kernel primarily leads to drops in the challenging contiguous and modulo schemes. This shows that the structure kernel is crucial for characterizing unseen sites in the wild type protein. While removing the site comparison and mutation probability kernels leads to small and medium drops in performance, we observe that removing both leads to an even larger performance drop, indicating a synergy between the two. In Table E.2, the ablation results per functional category are shown, where observe that the inclusion of the structure kernel is crucial for the increased performance in structure-related prediction tasks such as binding and stability.
@@ 4.2 Uncertainty quantification per mutation domain
Figure 2 shows the distributions of mean predictive variances in the six domains. In the three single mutant domains, we observe that the uncertainties increase from scheme to scheme, reflecting the difficulties of the tasks and analogously the expected performance scores (Table 1). When training on both single and double mutants (1M/2M), we observe a lower uncertainty on double mutants than single mutants. For many of the multi-mutant datasets, the mutants are not uniformly sampled but often include a fixed single mutation. A possible explanation is thus that it might be more challenging to decouple the signal from a double mutation into its constituent single mutation signals. In the extrapolation setting, we observe large predictive uncertainties, as expected. One explanation of the discrepancy between the variance distributions in the multi-mutant domains might lie in the difference in target distributions between training and test sets. Figure I.1 in the appendix shows the overall target distribution of assays for the 51 considered multi-mutant datasets. The single and double mutants generally belong to different modalities, where the double mutants often lead to a loss of fitness. This shows the difficulty of predicting on domains not encountered during training. For reference, we include the results for the multi-mutant domains in Table G.1 in the appendix.
4.3 Uncertainty calibration analysis
To clarify the relationship between model uncertainty and expected performance, we proceed with a calibration analysis. First, we perform a confidence interval-based calibration analysis [41], resulting in calibration curves which in the classification setting are known as reliability diagrams [68]. The results for each dataset are obtained via five-fold cross validation, corresponding to five separately trained models for each split scheme. We select four diverse datasets as examples (Table 3), reflecting both high and low predictive performance. The mean calibration curves can be seen in Figure 3(a). For method details and results across all datasets, see Appendix J. The mean expected calibration error (ECE) is shown in the bottom of each plot, where a value of zero indicates perfect calibration. Overall, the uncertainties appear to be well-calibrated both qualitatively from the curves and quantitatively from the ECEs. Even the smallest dataset (fourth row, $N=165$) achieves decent calibration, albeit with larger variances between folds.
@@ 4.3 Uncertainty calibration analysis
BLAT_ECOLX 0.804 0.826 0.909 0.846 4996 286 Organismal fitness [69]
PA_I34A1 0.226 0.457 0.539 0.407 1820 716 Organismal fitness [70]
TCRG1_MOUSE 0.849 0.849 0.928 0.875 621 37 Stability [4]
OPSD_HUMAN 0.739 0.734 0.727 0.734 165 348 Expression [71]
While the confidence interval-based analysis suggested that the uncertainty estimates are well-calibrated with some under-confidence, the same is not as visibly clear for the error-based calibration plots, suggesting that the expected correlation between model uncertainty and prediction error is not a given. We do however see an overall trend of increasing error with increasing uncertainty in three of the four datasets, where the curves lie close to the diagonal (as indicated by the dashed line). The second row shows poorer calibration – particularly in the modulo and contiguous schemes. The curves however remain difficult to interpret, in part due to the errorbars on both axes. To alleviate this, we compute similar metrics for all 217 datasets across the three splits, which indicate that, though varying, most calibration curves are well-behaved (see Section J.3).
We supplement the above calibration curves with Figures K.1, K.2, K.3 and K.4 in the Appendix, which show the true values plotted against the predictions. These highlight the importance of well-calibrated uncertainties and underline their role in interpreting model predictions and their trustworthiness.
@@ 4.3.1 Comparison with ProteinNPT
Monte Carlo (MC) dropout [48] is a popular uncertainty quantification technique for deep learning models. Calibration curves for ProteinNPT with MC dropout for the same four datasets across the three split schemes can be seen in Figure L.1 while figures showing the true values plotted against the predictions with uncertainties are shown in Section L.3. These indicate that employing MC dropout on a deep learning model like ProteinNPT seems to provide lower levels of calibration, providing overconfident uncertainties across assays and splits. Due to the generally low uncertainties and the resulting difference in scales, the calibration curves are often far from the diagonal. The trends in the calibration curves however show that the model errors often correlate with the uncertainties, suggesting that the model can be recalibrated to achieve decent calibration.
@@ 5 Discussion
We have shown that a carefully constructed Gaussian process is able to reach state-of-the-art performance for supervised protein variant effect prediction while providing reasonably well-calibrated uncertainty estimates. For a majority of datasets, this is achieved orders of magnitude faster than competing methods.
While the predictive performance on the substitution benchmark is an improvement over previous methods, our proposed model has its limitations. Due to the site-comparison mechanism, our model is unable to handle insertions and deletions as it only operates on a fixed structure. Additionally, as the number of mutations increases, the assumption of a fixed structure might worsen, depending on the introduced mutations, which can affect reliability as the local environments might change. An additional limitation is the GP’s $\mathcal{O}(N^{3})$ scaling with dataset size. While not a major obstacle in the single mutant setting, dataset sizes can quickly grow when handling multiple mutants. The last decades have however produced a substantial literature on algorithms for scaling GPs to larger datasets [72, 73, 74], which could alleviate the issue, and we therefore believe this to be a technical rather than fundamental limitation. An additional limitation might present itself it in the multi-mutant setting, where the lack of explicit modeling of epistasis can potentially hinder extrapolation to higher-order mutants, prompting further investigation.
@@
This work was funded in part by Innovation Fund Denmark (1044-00158A), the Novo Nordisk Synergy grant (NNF200C0063709), VILLUM FONDEN (40578), the Pioneer Centre for AI (DNRF grant number P1), and the Novo Nordisk Foundation through the MLSS Center (Basic Machine Learning Research in Life Science, NNF20OC0062606).
@@
Appendix A License and code availability
The codebase is publicly available at https://github.com/petergroth/kermut under the open source MIT License.
Appendix B GP details
B.1 Structure kernel
@@
The sum over all pairs of mutations in equation 2 is motivated by [63] who showed that a linear model was sufficient to effectively model the mutation effects. Note that $\text{Cov}(Y_{1}+Y_{2},Y_{3}+Y_{4})=\text{Cov}(Y_{1},Y_{3})+\text{Cov}(Y_{1},Y_{4})+\text{Cov}(Y_{2},Y_{3})+\text{Cov}(Y_{2},Y_{4})$. Hence, if the $Y_{i}$’s are the effects of single mutations, we can find the covariance between two double mutant variants by calculating the pairwise sum and similarly for other number of mutations.
B.2 Sequence kernel
@@ B.3 Parametrization note
B.4 Zero-shot mean function
@@ B.5 Kernel proof
B.6 Automatic Model Selection
Kermut has been developed from the ground up as described Section 3.3, which led to separate structure and sequence kernels which are added together. Alternatively, an automatic model selection scheme can be employed for data-driven kernel composition as proposed in Chapter 3 in [78]. For demonstrative purposes, we conduct such a model selection procedure. We define four base kernels ($k_{H}$, $k_{p}$, $k_{d}$, $k_{\text{seq}}$) as well as sum and product operations. We choose a subset of 17 ProteinGym assays and fit a GP (with a zero-shot mean function) and each of the four kernels, where the best performing kernel across splits is kept. For the second round, the remaining three kernels are either added or multiplied to the existing kernel. This process is continued until either all four kernels are used or until the test performance no longer increases. The results can be seen in Table B.1, where the final kernel is the product of the four base kernels. The results on the 174 ablation datasets using this kernel (denoted as "Kermut (product)") can be seen with the main Kermut GP in Table F.2. The ECE/ENCE values for Kermut (product) are 0.060/0.192 vs. 0.051/0.170 for the main formulation. Due to the potential for interpretability of the sum-formulation as well as slightly better calibration, the main formulation of the Kermut GP remains as in Equation 4.
Table B.1: Automatic Model Selection with four base kernels and sum/product operations. The resulting kernel is the product of all base kernels.
Kernel Round 1 Round 2 Round 3 Round 4
$k_{\text{seq}}$ 0.555
$k_{d}$ 0.589
$k_{H}$ 0.584
$k_{p}$ 0.521
$k_{d}\times k_{\text{seq}}$ 0.653
$k_{d}+k_{\text{seq}}$ 0.642
$k_{d}\times k_{H}$ 0.623
$k_{d}+k_{H}$ 0.622
$k_{d}\times k_{p}$ 0.633
$k_{d}+k_{p}$ 0.629
$k_{d}\times k_{\text{seq}}\times k_{H}$ 0.666
$k_{d}\times k_{\text{seq}}+k_{H}$ 0.657
$k_{d}\times k_{\text{seq}}\times k_{p}$ 0.678
$k_{d}\times k_{\text{seq}}+k_{p}$ 0.666
$k_{d}\times k_{\text{seq}}\times k_{p}\times k_{H}$ 0.684
$k_{d}\times k_{\text{seq}}\times k_{p}+k_{H}$ 0.676
Appendix C Implementation details
C.1 General details
We build our kernel using the GPyTorch framework [79]. We assume a homoschedastic Gaussian noise model, on which we place a HalfCauchy prior [80] with scale $0.1$. We fit the hyperparameters by maximizing the exact marginal likelihood with gradient descent using the AdamW optimizer [81] with learning rate 0.1 for a 150 steps, which proved to be sufficient for convergence for a number of sampled datasets.
C.2 System details
@@ C.3 Compute time
Scaling the experiments to the full ProteinGym benchmark is however costly. We were able to train/evaluate Kermut on 215/217 datasets using an NVIDIA A40 GPU with 48GB VRAM. The remaining two, POLG_CXB3N_Mattenberger_2021 and POLG_DEN26_Suphatrakul_2023, datasets were too large to fit into GPU memory without resorting to reduced precision. For these, we trained and evaluated the model using CPU only which takes considerable time.
C.4 Handling long sequences
C.4.1 Example wall-clock time
The wall-clock times for generating the results used for the calibration curves in Figures 3(b), 3(a) and L.1 for one split scheme can be seen in Table C.1 for Kermut and ProteinNPT. The experiments were carried out using identical hardware. The test system is however a shared compute cluster with sharded GPUs, so variance is expected between runs. Generating the full results for the figures thus takes Kermut approximately 10 minutes while ProteinNPT takes 400 hours. This shows the significant reduction in computational burden that Kermut allows for. Both ProteinNPT and Kermut assumes that sequence embeddings are available a priori.
Table C.1: Approximate wall clock times for training and evaluating Kermut and ProteinNPT for a single split scheme, i.e., by using 5-fold cross validation. While the runtime of Kermut scales with dataset size, ProteinNPT appears to scale more strongly with sequence length due to the tri-axial attention.